OpenStudy (anonymous):

write y-4=-3/2(x+6) in standard form

OpenStudy (igreen):

First distribute -3/2 into the parenthesis: y - 4 = -3/2(x + 6) y - 4 = -3/2x - 9 Add 4 to both sides, what's -9 + 4?

OpenStudy (igreen):

Can you add -9 + 4? @Oyin3

OpenStudy (anonymous):

@iGreen it is -5

OpenStudy (igreen):

Yep, so we have: y = -3/2x - 5 Now we add 3/2x to both sides: 3/2x + y = -5 That's our answer.

OpenStudy (igreen):

\(\bbox [10pt, lime, border:5pt solid black]{\huge\cal\color{red}\succ\color{blue}{Welcome\ to\ \color{#00A1FF}{Open}\color{#329932}{Study!\color{red}\prec}}}\) You can give medals by clicking 'Best Response'.